Harry has $4000 invested in 2 savings accounts. One account earns 6% interest per year, and the other pays 7% per year. If his total interest for the year is $264, how much is invested at 7%?---isnt it $167

Answer by 1guru(1)

Since we have to find the amount invested at 7%, Let this amount be y.
So, the amount invested at 6% is (4000-y)
7% interest on x is 7x/100 =.07y
similarly, 6% interest on (4000-y) is .06(4000-y)
This equals (.06x4000 + .06y)= 240 - .06y
The total of these interests is 264
So, .07y + 240 - .06y = 264
Or, .01y = 16
Or, y= 1600.
This is the amount invested at 7%
You can check this out:
7% x 1600 = 168
6% x (4000-1600) = 96
168+ 96 = 264
Note: The variable y (and not x) has been used to avoid confusion between the variable x and the multiplication sign x .